# Complexity of deep computations via topology of function spaces

### v3

@@ 3.2. The Todorčević trichotomy, and levels of NIP and PAC learnability
    Alexandroff Duplicate. Let $K$ be any compact metric space and consider the Polish space $X=K\sqcup C(K)$, i.e., the disjoint union of $C(K)$ (with its supremum norm topology) and $K$. For each $a\in K$ define $g_{a}^{0},g_{a}^{1}:X\rightarrow\mathbb{R}$ as follows: g_a^0(x)={0, if x∈Kx(a), if x∈C(K); g_a^1(x)={δa(x), if x∈K;x(a), if x∈C(K). Let $D(K)=\{g_{a}^{0}:a\in K\}\cup\{g_{a}^{1}:a\in K\}$. Observe that all points $g_{a}^{1}$ are isolated and that open neighborhoods of $g_{a_{0}}^{0}$ are of the form $g_{a}^{i}:a\in U,i\in\{0,1\}\}\backslash\{g_{a}^{1}:a\in F\}$ where $U\subseteq K$ is an open neighborhood of $a_{0}$ and $F\subseteq K$ is a finite set. Another abstract way in which this space is presented is as the space $K\times\{0,1\}$ whose basic open neighborhoods are given as before, identifying $(a,0)\mapsto g_{a}^{0}$ and $(a,1)\mapsto g_{a}^{1}$. We can also embed $D(K)$ into the product $A(K)\times K$ by identifying $(a,0)\mapsto(\mathbf{0},a)$ and $(a,1)\mapsto(\delta_{a},a)$. Notice that $D(K)$ is a first countable Rosenthal compactum. It is not separable if $K$ is uncountable, thus we typically study the interesting case when $K=2^{\mathbb{N}}$. As with the Alexandroff compactification $A(2^{\mathbb{N}})$, we can make the space $D(2^{\mathbb{N}})$ separable by adding a countable set. For example, the closure of the set $\{(v_{s},s^{\frown}0^{\infty}):s\in 2^{<\mathbb{N}}\}\subseteq\hat{A}(2^{\mathbb{N}})\times 2^{\mathbb{N}}$ is $\{(\mathbf{0},a):a\in 2^{\mathbb{N}}\}\cup\{(\delta_{a},a):a\in 2^{\mathbb{N}}\}\cup\{(v_{s},s^{\frown}0^{\infty}):s\in 2^{<\mathbb{N}}\}$, where $\{(\mathbf{0},a):a\in 2^{\mathbb{N}}\}\cup\{(\delta_{a},a):a\in 2^{\mathbb{N}}\}$ is homeomorphic to $D(2^{\mathbb{N}})$.
@@
    Extended Alexandroff Duplicate of the Split Cantor. For each finite binary sequence $t\in 2^{<\mathbb{N}}$ let $a_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with 0’s and let $b_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with $1$’s. Define $h_{t}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ by h_t(x)={0, if x¡at;1/2, if at≤x ≤bt;1, if bt¡x. Let $\hat{D}(S(2^{\mathbb{N}}))$ be the pointwise closure of the set $\{h_{t}:t\in 2^{<\mathbb{N}}\}$. The identification $h_{t}\mapsto(v_{t},f_{t^{\frown}0^{\infty}}^{+})$ lifts to a homeomorphism between $\hat{D}(S(2^{\mathbb{N}}))$ and the subspace of $\hat{A}(2^{\mathbb{N}})\times S(2^{\mathbb{N}})$ consisting of $(\mathbf{0},f_{a}^{+})$, $(\mathbf{0},f_{a}^{-})$, $(\delta_{a},f_{a}^{+})$ and $(v_{t},f_{t^{\frown}0^{\infty}}^{+})$ for $a\in 2^{\mathbb{N}}$ and $t\in 2^{<\mathbb{N}}$ (see 4.3.7 in [ADK08]). Hence, $\hat{D}(S(2^{\mathbb{N}}))$ is a separable first countable Rosenthal compactum which is not hereditarily separable. In fact, it contains an uncountable discrete subspace.
@@
    Let $(L,\mathcal{P},\Gamma)$ be the CCS given by $L=2^{\mathbb{N}}$, $\mathcal{P}=\{P_{n}:n\in\mathbb{N}\}$ and $\Gamma$ is the semigroup generated by $\Delta=\{\gamma_{t}:t\in 2^{<\mathbb{N}}\}$, where $P_{n}:2^{\mathbb{N}}\rightarrow\{0,1\}$ is the projection map $P_{n}(x)=x(n)$ and $\gamma_{t}:L\rightarrow L$ is given by γ_t(x)={0, if lext0;(01), if t0lexx ≤lext1;1, if t1¡lexx. where $(01)^{\infty}$ denotes the sequence of alternating bits: $010101\cdots$. As in the other examples, it is not difficult to see that $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom. For example, the condition $t^{\frown}0^{\infty}\leq_{\text{lex}}x\leq_{\text{lex}}t^{\frown}1^{\infty}$ is equivalent to $x$ extending $t$. Observe that the set of deep computations is homeomorphic to $\hat{D}(S(2^{\mathbb{N}}))$ (see Example 3.3(5)). This is an example of $\Delta$ which is NIP1 but not NIP2.
@@ 4. Randomized versions of NIP and Monte Carlo computability of deep computations
In this section, we replace deterministic computability by probabilistic (‘Monte Carlo’) computability. We do not assume that $\mathcal{P}$ is countable. The main results of the section are Theorem 4.8 (connecting NIP and Monte Carlo computability) and 4.14 (connecting Talagrand stability and Monte Carlo computability).
@@ 4.1. NIP and Monte Carlo computability of deep computations
The raison d’être of the Baire class-1 functions is to have with a class of functions that are obtained as equential limit points of continuous functions. By Fact 1.2, for perfectly normal $X$, a function $f$ is in $B_{1}(X,Y)$ if and only if $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ for every open $U\subseteq Y$. Thus, for such $X$, functions in $B_{1}(X,Y)$ are not too far from being continuous. In this section we will study a more general class of functions, namely, the class of universally measurable functions, which we define next.
@@
In the context of deep computations, we are interested in transition maps of a state space $L\subseteq\mathbb{R}^{\mathcal{P}}$ into itself. There are two natural $\sigma$-algebras one can consider in the product space $\mathbb{R}^{\mathcal{P}}$: the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by open sets in $\mathbb{R}^{\mathcal{P}}$, and the cylinder $\sigma$-algebra (i.e., the $\sigma$-algebra generated by the sub-basic open sets in $\mathbb{R}^{\mathcal{P}}$, that is, the sets $\pi_{P}^{-1}(U)$ with $U\subseteq\mathbb{R}$ open and $P\in\mathcal{P}$). Note that when $\mathcal{P}$ is countable, both $\sigma$-algebras coincide, but in general the cylinder $\sigma$-algebra is strictly smaller. We will use the cylinder $\sigma$-algebra to define universally measurable maps $f:\mathbb{R}^{\mathcal{P}}\rightarrow\mathbb{R}^{\mathcal{P}}$. The reason for this choice is the following characterization:
@@
The preceding proposition says that a transition map is universally measurable if and only if it is universally measurable on all its features; in other words, we can check measurability of a transition just by checking measurability feature by feature. This is the same as in the Baire class-1 case (compare with Proposition 1.10).
The main result in section 3 is that, as long as we work with countably many features, PAC-learning (or NIP) corresponds to relative compactness in the space of Baire class-1 functions. The following result (which does not assume countability of the number of features) gives an analogous characterization of the NIP in terms of universal measurability:
@@
Condition (2) of Theorem 4.5 shows that to study measure-theoretic versions of NIP, we need only consider compact subsets of $X$. Now, every Radon measure on a compact space is finite; hence, by proper normalization, it can be treated as a probability measure. Therefore, in the context of Monte Carlo measurability, we focus on Radon probability measures rather than general Radon measures.
@@ 4.2. Talagrand stability and Monte Carlo computability of deep computations
    $(X,\Sigma,\mu)$ is inner regular with respect to $\mathcal{K}$, i.e., for all $E\in\Sigma$ μ(E)=sup{μ(K):$K$ and KE}.
@@
Notice that the preceding two results apply to sets of measurable functions, a class of functions larger than the class of continuous functions. However, by the Extendibility Axiom, finitary computations are continuous, i.e., if $A$ is a set of computations, then $A\subseteq C_{p}(X)$. The question of whether we can remove the set-theoretic assumption in Theorem 4.15 when $A\subseteq C_{p}(X)$ (instead of $A\subseteq M_{r}(X)$) remains open.

### v4

@@ 3.2. The Todorčević trichotomy, and levels of NIP and PAC learnability
    Alexandroff Duplicate. Let $K$ be any compact metric space and consider the Polish space $X=K\sqcup C(K)$, i.e., the disjoint union of $C(K)$ (continuous functions on $K$ with the supremum norm topology) and $K$. For each $a\in K$ define $g_{a}^{0},g_{a}^{1}:X\rightarrow\mathbb{R}$ as follows:
    $g_{a}^{0}(x)=\begin{cases}0,&\text{if }x\in K\\ x(a),&\text{if }x\in C(K);\end{cases}$
    $g_{a}^{1}(x)=\begin{cases}\delta_{a}(x),&\text{if }x\in K;\\ x(a),&\text{if }x\in C(K).\end{cases}$
    Let $D(K)=\{g_{a}^{0}:a\in K\}\cup\{g_{a}^{1}:a\in K\}$. Observe that all points $g_{a}^{1}$ are isolated and that open neighborhoods of $g_{a_{0}}^{0}$ are of the form $\{g_{a}^{i}:a\in U,i\in\{0,1\}\}\setminus\{g_{a}^{1}:a\in F\}$ where $U\subseteq K$ is an open neighborhood of $a_{0}$ and $F\subseteq K$ is a finite set. Another abstract way in which this space is presented is as the space $K\times\{0,1\}$ whose basic open neighborhoods are given as before, identifying $(a,0)\mapsto g_{a}^{0}$ and $(a,1)\mapsto g_{a}^{1}$. We can also embed $D(K)$ into the product $A(K)\times K$ by identifying $(a,0)\mapsto(\mathbf{0},a)$ and $(a,1)\mapsto(\delta_{a},a)$. Notice that $D(K)$ is a first countable Rosenthal compactum. It is not separable if $K$ is uncountable; thus, we typically study the interesting case when $K=2^{\mathbb{N}}$. As with the Alexandroff compactification $A(2^{\mathbb{N}})$, we can make the space $D(2^{\mathbb{N}})$ separable by adding a countable set. For example, the closure of the set $\{(v_{s},s^{\frown}0^{\infty}):s\in 2^{<\mathbb{N}}\}\subseteq\hat{A}(2^{\mathbb{N}})\times 2^{\mathbb{N}}$ is $\{(\mathbf{0},a):a\in 2^{\mathbb{N}}\}\cup\{(\delta_{a},a):a\in 2^{\mathbb{N}}\}\cup\{(v_{s},s^{\frown}0^{\infty}):s\in 2^{<\mathbb{N}}\}$, where $\{(\mathbf{0},a):a\in 2^{\mathbb{N}}\}\cup\{(\delta_{a},a):a\in 2^{\mathbb{N}}\}$ is homeomorphic to $D(2^{\mathbb{N}})$.
@@
    Extended Alexandroff Duplicate of the Split Cantor. For each finite binary sequence $t\in 2^{<\mathbb{N}}$ let $a_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with 0’s and let $b_{t}\in 2^{\mathbb{N}}$ be the sequence starting with $t$ and ending with $1$’s. Define $h_{t}:2^{\mathbb{N}}\rightarrow\mathbb{R}$ by
    $h_{t}(x)=\begin{cases}0,&\text{if }x<a_{t};\\ 1/2,&\text{if }a_{t}\leq x\leq b_{t};\\ 1,&\text{if }b_{t}<x.\end{cases}$
    Let $\hat{D}(S(2^{\mathbb{N}}))$ be the pointwise closure of the set $\{h_{t}:t\in 2^{<\mathbb{N}}\}$. The identification $h_{t}\mapsto(v_{t},f_{t^{\frown}0^{\infty}}^{+})$ lifts to a homeomorphism between $\hat{D}(S(2^{\mathbb{N}}))$ and the subspace of $\hat{A}(2^{\mathbb{N}})\times S(2^{\mathbb{N}})$ consisting of $(\mathbf{0},f_{a}^{+})$, $(\mathbf{0},f_{a}^{-})$, $(\delta_{a},f_{a}^{+})$ and $(v_{t},f_{t^{\frown}0^{\infty}}^{+})$ for $a\in 2^{\mathbb{N}}$ and $t\in 2^{<\mathbb{N}}$ (see 4.3.7 in [ADK08]). Hence, $\hat{D}(S(2^{\mathbb{N}}))$ is a separable first countable Rosenthal compactum which is not hereditarily separable. In fact, it contains an uncountable discrete subspace.
@@
    Let $(L,\mathcal{P},\Gamma)$ be the CCS given by $L=2^{\mathbb{N}}$, $\mathcal{P}=\{P_{n}:n\in\mathbb{N}\}$ and $\Gamma$ is the semigroup generated by $\Delta=\{\gamma_{t}:t\in 2^{<\mathbb{N}}\}$, where $P_{n}:2^{\mathbb{N}}\rightarrow\{0,1\}$ is the projection map $P_{n}(x)=x(n)$ and $\gamma_{t}:L\rightarrow L$ is given by
    $\gamma_{t}(x)=\begin{cases}0^{\infty},&\text{if }x<_{\text{lex}}t^{\frown}0^{\infty};\\ (01)^{\infty},&\text{if }t^{\frown}0^{\infty}\leq_{\text{lex}}x\leq_{\text{lex}}t^{\frown}1^{\infty};\\ 1^{\infty},&\text{if }t^{\frown}1^{\infty}<_{\text{lex}}x.\end{cases}$
    where $(01)^{\infty}$ denotes the sequence of alternating bits: $010101\cdots$. As in the other examples, it is not difficult to see that $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom. For example, the condition $t^{\frown}0^{\infty}\leq_{\text{lex}}x\leq_{\text{lex}}t^{\frown}1^{\infty}$ is equivalent to $x$ extending $t$. Observe that the set of deep computations is homeomorphic to $\hat{D}(S(2^{\mathbb{N}}))$ (see Example 3.3(5)). This is an example of $\Delta$ which is NIP1 but not NIP2.
@@ 4. Randomized versions of NIP and Monte Carlo computability of deep computations
In this section, we replace deterministic computability by probabilistic (‘Monte Carlo’) computability. We do not assume that $\mathcal{P}$ is countable. (The countability assumption on $\mathcal{P}$ played a crucial role in the proof of Theorem 1.11, as it makes $\mathbb{R}^{\mathcal{P}}$ a Polish space.) The main results of the section are Theorem 4.8 (connecting NIP and Monte Carlo computability) and 4.14 (connecting Talagrand stability and Monte Carlo computability).
@@ 4.1. NIP and Monte Carlo computability of deep computations
The raison d’être of the Baire class 1 functions is to have a class of functions that are obtained as sequential limit points of continuous functions. By Fact 1.2, for perfectly normal $X$, a function $f$ is in $B_{1}(X,Y)$ if and only if $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ for every open $U\subseteq Y$. Thus, for such $X$, functions in $B_{1}(X,Y)$ are not too far from being continuous. In this section we will study a more general class of functions, namely, the class of universally measurable functions, which we define next.
@@
In the context of deep computations, we are interested in transition maps of a state space $L\subseteq\mathbb{R}^{\mathcal{P}}$ into itself. In the product space $\mathbb{R}^{\mathcal{P}}$, we have two natural $\sigma$-algebras: the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by open sets in $\mathbb{R}^{\mathcal{P}}$, and the cylinder $\sigma$-algebra (i.e., the $\sigma$-algebra generated by the sub-basic open sets in $\mathbb{R}^{\mathcal{P}}$, that is, the set $\pi_{i}^{-1}(U)$ with $U\subseteq\mathbb{R}$ open). In general, the cylinder $\sigma$-algebra is strictly smaller, but when $\mathcal{P}$ is countable, both $\sigma$-algebras coincide. We will use the cylinder $\sigma$-algebra to define universally measurable maps $f:\mathbb{R}^{\mathcal{P}}\rightarrow\mathbb{R}^{\mathcal{P}}$. The reason for this choice is the following characterization:
@@
The preceding proposition says that a transition map is universally measurable if and only if it is universally measurable on all its features; in other words, we can check measurability of a transition just by checking measurability feature by feature. This is the same as in the Baire class 1 case (compare with Proposition 1.10).
The main result in section 3 is that, as long as we work with countably many features, PAC-learning (or NIP) corresponds to relative compactness in the space of Baire class 1 functions. The following result (which does not assume countability of the number of features) gives an analogous characterization of the NIP in terms of universal measurability:
@@
Every Radon measure on a compact space (e.g., $\mathcal{L}_{\text{sh}}$) is finite; hence, by proper normalization, it can be treated as a probability measure. For this reason, in the context of Monte Carlo computability (4.6), we restrict our attention to Radon probability measures instead of general Radon measures.
@@ 4.2. Talagrand stability and Monte Carlo computability of deep computations
    $(X,\Sigma,\mu)$ is inner regular with respect to $\mathcal{K}$, i.e., for all $E\in\Sigma$
    $\mu(E)=\sup\{\mu(K):K\in\mathcal{K}\text{ and }K\subseteq E\}.$
@@
Notice that the preceding two results apply to sets of measurable functions, a class of functions larger than the class of continuous functions. However, by the Extendibility Axiom, finitary computations are continuous, i.e., if $A$ is a set of computations, then $A\subseteq C_{p}(X)$.
5. Conclusion
We have seen how the tools of model theory and the topology of function spaces can give us new insights into classifying the difficulty of computations. In future papers, we plan to extend this research in a variety of directions. A direction of particular interest to us is replacing classical computation by quantum computation. Topologically, our results so far are not optimal because of their restriction to Polish spaces. We expect to extend most of them to $K$-countably determined (Lindelöf $\Sigma$) spaces or $K$-analytic spaces, which have previously been used in functional analysis — see e.g., Kakol et al [KKL+02]. Our restricting our considerations to continuous and Baire class 1 functions can be relaxed by applying more high-powered model-theoretic and topological techniques. This is work in progress. We have hinted at applications to deep equilibria; see [ADI+24]. In another direction, we have hinted at the role additional set-theoretic axioms may play. Since real world computations are finite and hence do not depend on other than the usual set-theoretic axioms, in principle we can assume these axioms and then dispense with them; however, this is delicate, since many of our techniques involve infinite sets so we have to check that our assertions are sufficiently simple so as to be absolute, i.e., unaffected by additional axioms. We hope that this paper, the preceding one [ADI+24], and the future ones will stimulate interactions among these three seemingly unrelated fields of research.